The equilibrium constant, Kc, for the following reaction is 1.31×10-2 at 700 K.


NH4I(s) NH3(g) + HI(g)



Calculate Kc at this temperature for the following reaction:


NH3(g) + HI(g) NH4I(s)


Kc = 


1
Expert's answer
2022-04-20T09:20:04-0400

NH4I(s) = NH3(g) + HI(g), Kc = 1.31×10-2

NH3(g) + HI(g) = NH4I(s), Kc' = 1/Kc = 1 / 1.31×10-2 = 76.34